OpenStudy (anonymous):

simplify : ∛(22√5 + 15√11) + ∛(22√5 - 15√11)

OpenStudy (anonymous):

give me a clue for do this??

OpenStudy (lgbasallote):

honestly..you cant do anything about this...

OpenStudy (anonymous):

i was also thinking the same..

OpenStudy (anonymous):

use \[(a+b)^{3} = a ^{3}+b ^{3}+3\times a \times b(a+b)\] Put a+b = t where a+b is ur question. solve the cubic.

OpenStudy (anonymous):

sorri, the problem should be ∛(22√5 + 15√11) + ∛(22√5 - 15√11) - \[\sqrt[6]{3520}\], i want do it step by step only, hehe ... can u exlaind to me?

OpenStudy (anonymous):

\[\huge \sqrt[?]{3520} ?\]

OpenStudy (anonymous):

? is 6

OpenStudy (anonymous):

let a=22√5 + 15√11 and b=22√5 - 15√11 ab=-55=-5*11

OpenStudy (anonymous):

there is something wrong with equation editor let me reload the page

OpenStudy (anonymous):

\[\Large 22\sqrt{5} = 2\sqrt{11}\sqrt{55}\] \[\Large15\sqrt{11} = 3\sqrt{5}\sqrt{55}\] \[\Large \sqrt[6]{3520} = 2\sqrt[6]{55} = 2\sqrt[3]{\sqrt{55}}\] \[\Large \sqrt[3]{(22\sqrt{5}+15\sqrt{11})}+\sqrt[3]{(22\sqrt{5}-15\sqrt{11})}-\sqrt[6]{3520}\] \[\Large \sqrt[3]{(2\sqrt{11}\sqrt{55}+3\sqrt{5}\sqrt{55})}+\sqrt[3]{(2\sqrt{11}\sqrt{55}-3\sqrt{5}\sqrt{55})}-2\sqrt[3]{\sqrt{55}}\] \[\Large A = \sqrt{55}\] \[\Large \sqrt[3]{(2\sqrt{11}A+3\sqrt{5}A)}+\sqrt[3]{(2\sqrt{11}A-3\sqrt{5}A)}-2\sqrt[3]{A}\] \[\Large \sqrt[3]{A(2\sqrt{11}+3\sqrt{5})}+\sqrt[3]{A(2\sqrt{11}-3\sqrt{5})}-2\sqrt[3]{A}\] \[\Large \sqrt[3]{A}\sqrt[3]{(2\sqrt{11}+3\sqrt{5})}+\sqrt[3]{A}\sqrt[3]{(2\sqrt{11}-3\sqrt{5})}-2\sqrt[3]{A}\] \[\Large \sqrt[3]{A}(\sqrt[3]{(2\sqrt{11}+3\sqrt{5})}+\sqrt[3]{(2\sqrt{11}-3\sqrt{5})}-2)\] Another way to look at this: \[\Large B = 2\sqrt{11} \] \[\Large C = 3\sqrt{5}\] \[\Large \sqrt[3]{A}(\sqrt[3]{(B+C)}+\sqrt[3]{(B-C)}-2)\] That's about as far as I got. Can't recall if the sum and difference of cubes equations can be used on cube roots.